OpenStudy (anonymous):

f(2) = −4, g(2) = 2, f '(2) = −5 and g'(2) = 3. Find h'(2). h(x) = g(x) / 1 + f(x)

OpenStudy (kirbykirby):

You could use the quotient rule. Let's just say \(\color{red}{g(x)=u(x)}\) \(\color{blue}{1+f(x)=w(x)}\) I'm just using this notation so you can see the quotient rule. \[ h(x)=\frac{\color{red}{u}}{\color{blue}{w}}\], so by the quotient rule: \[ h'(x)=\frac{\color{red}{u}'\color{blue}{w}-\color{red}{u}\color{blue}{w}'}{\color{blue}{w}^2}\].. substituting what you actually have: \[\begin{align}h'(x)&=\frac{(\color{red}{g(x)})'\cdot[\color{blue}{1+f(x)}]-\color{red}{g(x)}\cdot[\color{blue}{1+f(x)}]'}{[\color{blue}{1+f(x)}]^2}\\~\\&=\frac{g'(x)[1+f(x)]-g(x)[f'(x)]}{[1+f(x)]^2}\end{align}\] so: \[ h'(2)=\frac{g'(2)[1+f(2)]-g(2)[f'(2)]}{[1+f(2)]^2}\]